Use the method of cylinders to determine the volume of the solid obtained by rotating the region bounded by y=1/x, x=1/2, x=4 and the x-axis about the y-axis.


1
Expert's answer
2021-05-02T08:11:42-0400

The volume of the solid generated by rotating this region about the y- axis may be found by adding the volumes of the solids generated by rotating the region bounded by "x=4", "x= \frac 12", "y=\frac1x" (v1) and the region bounded by "x=4", "x=\frac 12", "y=1\/4" and the x-axis (v2).

"v_1=2π∫_0^{\frac14}yx dy"

"=2π∫_0^{\frac14}y(4-\frac 12) dy"

"=2π∫_0^{\frac14)} {\frac72} y dy"

"= [ \frac {7π y^2}2]^{\frac14}_0"

"=7π [\frac{(\frac14)^2}2 -0]"

"=\frac 7{32}π cubic units of length"

"v_2=2π ∫_{\frac14}^2yx dy=2π∫_{\frac14}^2y(\frac1y- \frac 12)dy"

"=2π∫_{\frac14}^2(1- \frac12 y)dy"

"=2π[y-\frac14 y^2 ] ^2_{\frac14}"

"=2π[(2-\frac 14 (2)^2 )-(\frac 14-\frac 14 (\frac14)^2)]"

"=2π(1- \frac {15}{64})"

"=\frac {49}{32} π cubic units of length"

"The total volume V=v_1+v_2"

"=\frac7{32}π+ \frac {49}{32} π"

"=\frac 74 π cubic units of length"